OpenStudy (anonymous):

please help simplify.. (4x^3)^2 (2x^-4)

OpenStudy (anonymous):

\[\huge (Cx^a)^b=C^bx^{ab}\]

OpenStudy (anonymous):

\[\huge {x^a \over x^b}=x^{a-b}\]

OpenStudy (anonymous):

\[(4x^3)^2 (2x^{-4})={(4x^3)^2 (2) \over x^4}\]

OpenStudy (anonymous):

now use the two formulas i gave u to simplify it